# Equal contribution    Project lead    Corresponding author

### v2

@@ D.1 Ablation Study
Table 11: MSE for horizon-96 forecasting across six benchmarks, evaluated with different model components.
ETTh1 ETTh2 ETTm1 ETTm2 Weather Global Temp Average
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.211 0.262
         w/o Huber loss 0.365 0.309 0.344 0.205 0.163 0.217 0.267
         w/o multi-resolution layer 0.358 0.313 0.348 0.212 0.164 0.217 0.269
         w/o mixture-of-experts 0.370 0.317 0.347 0.212 0.163 0.223 0.272
         w/o auxiliary loss 0.368 0.325 0.350 0.219 0.164 0.226 0.275
As shown in Table 11, replacing the MoE layers with standard FFNs (denoted as “w/o mixture-of-experts ”) led to a noticeable performance decline, with the MSE worsening from $0.262$ to $0.272$. This highlights the significant contribution of the sparse architecture to the model’s overall performance, as its dynamic routing enables more specialized processing of diverse input patterns.
@@
Time-MoEbase 0.357 0.305 0.338 0.201 0.160 0.211 0.262 0.095 s/iter
@@ Appendix E Forecast Showcases
To visualize the performance differences among various large-scale time series models, we present the forecasting results of our model, Time-MoE, in comparison to the ground truth across six real-world benchmarks. These benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Global Temp datasets. Alongside Time-MoE’s results, we also show the performance of other large-scale baseline models at different scales, providing a comprehensive view of their comparative capabilities (Figures 510). In all figures, the context length is set to 512, and the forecast horizon is 96. To enhance clarity and aesthetics, we display the full forecast output, complemented by a portion of the preceding historical input data, ensuring a more intuitive comparison.
@@
Figure 6: Zero-shot forecasting cases from ETTh2 by different models, with forecast horizon 96.
Figure 7: Zero-shot forecasting cases from ETTm1 by different models, with forecast horizon 96.
Figure 8: Zero-shot forecasting cases from ETTm2 by different models, with forecast horizon 96.
Figure 9: Zero-shot forecasting cases from Weather by different models, with forecast horizon 96.
Figure 10: Zero-shot forecasting cases from Global Temp by different models, with forecast horizon 96.

### v3

@@ D.1 Ablation Study
Table 11: MSE and MAE for horizon-96 forecasting across six benchmarks, evaluated with different model components.
ETTh1 ETTh2 ETTm1 ETTm2 Weather Global Temp
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Time-MoEbase 0.357 0.381 0.305 0.359 0.338 0.368 0.201 0.291 0.160 0.214 0.211 0.343
         w/o Huber loss 0.365 0.383 0.309 0.366 0.344 0.369 0.205 0.295 0.163 0.221 0.217 0.359
         w/o multi-resolution layer 0.358 0.379 0.313 0.362 0.348 0.377 0.212 0.301 0.164 0.219 0.217 0.354
         w/o mixture-of-experts 0.370 0.398 0.317 0.372 0.347 0.373 0.212 0.298 0.163 0.218 0.223 0.357
         w/o auxiliary loss 0.368 0.394 0.325 0.387 0.350 0.377 0.219 0.304 0.164 0.220 0.226 0.363
As shown in Table 11, replacing the MoE layers with standard FFNs (denoted as “w/o mixture-of-experts ”) led to a noticeable performance decline, with the average MSE worsening from $0.262$ to $0.272$. This highlights the significant contribution of the sparse architecture to the model’s overall performance, as its dynamic routing enables more specialized processing of diverse input patterns.
@@
Time-MoEbase w/ {1,8,32,64} 0.357 0.305 0.338 0.201 0.160 0.211 0.262 0.095 s/iter
@@ D.2 Training Precision Analysis
D.3 Additional Experimental Results
D.3.1 TaxiBJ Dataset
We include a benchmark dataset, TaxiBJ (Zhang et al., 2017) for short-term forecasting evaluation. This original dataset encompasses taxicab GPS data and meteorological information collected from Beijing over four distinct intervals: July 1, 2013 - October 30, 2013; March 1, 2014 - June 30, 2014; March 1, 2015 - June 30, 2015; and November 1, 2015 - April 10, 2016. We selected the in-flow data from the period November 1, 2015, to April 10, 2016 as our benchmark. This benchmark dataset consists of 1,024 time-series sequences derived from $32\times 32$ grid cells.
We conducted evaluations on all zero-shot models using this benchmark, and set the context length to $512$ for all baselines. The results are summarized in Table 14.
Table 14: Short-term zero-shot forecasting results in TaxiBJ: A lower MSE or MAE indicates a better prediction. Red: the best, Blue: the 2nd best.
Metrics
TaxiBJ 1 0.214 0.294 0.214 0.292 0.214 0.294 0.334 0.373 0.282 0.334 0.267 0.323 0.247 0.316 0.866 0.751 0.250 0.315 0.255 0.316 0.250 0.303
8 0.302 0.363 0.297 0.356 0.302 0.362 0.487 0.470 0.427 0.422 0.431 0.425 0.393 0.430 0.883 0.759 0.341 0.380 0.311 0.352 0.310 0.351
24 0.385 0.419 0.376 0.410 0.385 0.417 0.610 0.529 0.530 0.477 0.548 0.488 0.494 0.495 0.894 0.764 0.438 0.440 0.427 0.420 0.431 0.418
48 0.423 0.448 0.414 0.440 0.422 0.444 0.626 0.542 0.559 0.497 0.563 0.500 0.524 0.515 0.892 0.765 0.502 0.478 0.475 0.450 0.494 0.460
The results indicate that our models consistently outperform other baselines in short-term forecasting on the TaxiBJ dataset.
D.3.2 Comparison to Timer, TFT, and N-BEATS
In this section, we incorporate additional baseline models for a more comprehensive evaluation. Specifically, Timer (2024d) is included for zero-shot forecasting (Table 15), while TFT (2021) and N-BEATS (2020) are included for in-domain forecasting (Table 16). The results indicate that our models consistently demonstrate improved performance relative to these established approaches.
Table 15: Additional zero-shot forecasting results of Timer, with 1B, 16B and 28B representing the scale of pretraining datasets as presented in the original paper Liu et al. (2024d). A lower MSE or MAE indicates a better prediction. Red: the best, Blue: the 2nd best.
Models [Uncaptioned image] Time-MoE (Ours) Timer
Time-MoEbase Time-MoElarge Time-MoEultra 1B 16B 28B
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.357 0.381 0.350 0.382 0.349 0.379 0.438 0.425 0.364 0.388 0.393 0.421
192 0.384 0.404 0.388 0.412 0.395 0.413 0.509 0.459 0.401 0.410 0.434 0.447
336 0.411 0.434 0.411 0.430 0.447 0.453 0.554 0.482 0.423 0.422 0.460 0.464
720 0.449 0.477 0.427 0.455 0.457 0.462 0.706 0.544 0.436 0.444 0.487 0.494
Avg. 0.400 0.424 0.394 0.419 0.412 0.426 0.552 0.478 0.406 0.416 0.444 0.456
ETTh2 96 0.305 0.359 0.302 0.354 0.292 0.352 0.315 0.351 0.294 0.350 0.308 0.369
192 0.351 0.386 0.364 0.385 0.347 0.379 0.393 0.402 0.353 0.385 0.348 0.398
336 0.391 0.418 0.417 0.425 0.406 0.419 0.412 0.422 0.376 0.400 0.366 0.414
720 0.419 0.454 0.537 0.496 0.439 0.447 0.425 0.440 0.393 0.420 0.409 0.446
Avg. 0.366 0.404 0.405 0.415 0.371 0.399 0.386 0.404 0.354 0.389 0.358 0.407
ETTm1 96 0.338 0.368 0.309 0.357 0.281 0.341 0.690 0.526 0.766 0.549 0.420 0.418
192 0.353 0.388 0.346 0.381 0.305 0.358 0.757 0.560 0.755 0.553 0.467 0.445
336 0.381 0.413 0.373 0.408 0.369 0.395 0.832 0.594 0.765 0.561 0.502 0.467
720 0.504 0.493 0.475 0.477 0.469 0.472 0.883 0.627 0.752 0.565 0.558 0.499
Avg. 0.394 0.415 0.376 0.405 0.356 0.391 0.791 0.577 0.760 0.557 0.487 0.457
ETTm2 96 0.201 0.291 0.197 0.286 0.198 0.288 0.213 0.295 0.234 0.312 0.247 0.324
192 0.258 0.334 0.250 0.322 0.235 0.312 0.283 0.339 0.287 0.343 0.294 0.358
336 0.324 0.373 0.337 0.375 0.293 0.348 0.346 0.377 0.340 0.373 0.335 0.385
720 0.488 0.464 0.480 0.461 0.427 0.428 0.424 0.424 0.437 0.426 0.386 0.418
Avg. 0.317 0.365 0.316 0.361 0.288 0.344 0.317 0.359 0.324 0.364 0.316 0.371
Weather 96 0.160 0.214 0.159 0.213 0.157 0.211 0.181 0.232 0.203 0.255 0.243 0.283
192 0.210 0.260 0.215 0.266 0.208 0.256 0.234 0.284 0.254 0.296 0.288 0.320
336 0.274 0.309 0.291 0.322 0.255 0.290 0.297 0.332 0.313 0.336 0.323 0.345
720 0.418 0.405 0.415 0.400 0.405 0.397 0.364 0.380 0.408 0.395 0.362 0.374
Avg. 0.265 0.297 0.270 0.300 0.256 0.288 0.269 0.307 0.294 0.321 0.304 0.331
Global Temp 96 0.211 0.343 0.210 0.342 0.214 0.345 0.250 0.373 0.245 0.372 0.308 0.425
192 0.257 0.386 0.254 0.385 0.246 0.379 0.299 0.415 0.300 0.418 0.359 0.465
336 0.281 0.405 0.267 0.395 0.266 0.398 0.347 0.451 0.365 0.466 0.415 0.507
720 0.354 0.465 0.289 0.420 0.288 0.421 0.452 0.521 0.542 0.585 0.579 0.617
Avg. 0.275 0.400 0.255 0.385 0.253 0.385 0.337 0.440 0.363 0.460 0.415 0.504
Average
0.336 0.384 0.336 0.380 0.322 0.372 0.394 0.406 0.378 0.393 0.344 0.391
Table 16: Additional results of in-domain forecasting baselines. A lower MSE or MAE indicates a better prediction. Red: the best, Blue: the 2nd best.
Time-MoEbase Time-MoElarge Time-MoEultra TFT N-BEATS
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.345 0.373 0.335 0.371 0.323 0.365 0.478 0.476 0.383 0.405
192 0.372 0.396 0.374 0.400 0.359 0.391 0.510 0.486 0.453 0.447
336 0.389 0.412 0.390 0.412 0.388 0.418 0.548 0.505 0.517 0.493
720 0.410 0.443 0.402 0.433 0.425 0.450 0.549 0.515 0.594 0.546
Avg. 0.379 0.406 0.375 0.404 0.373 0.406 0.521 0.496 0.487 0.473
ETTh2 96 0.276 0.340 0.278 0.335 0.274 0.338 0.352 0.387 0.362 0.384
192 0.331 0.371 0.345 0.373 0.330 0.370 0.429 0.432 0.413 0.430
336 0.373 0.402 0.384 0.402 0.362 0.396 0.461 0.460 0.430 0.448
720 0.404 0.431 0.437 0.437 0.370 0.417 0.475 0.473 0.554 0.530
Avg. 0.346 0.386 0.361 0.386 0.334 0.380 0.429 0.438 0.440 0.448
ETTm1 96 0.286 0.334 0.264 0.325 0.256 0.323 0.468 0.444 0.334 0.372
192 0.307 0.358 0.295 0.350 0.281 0.343 0.557 0.488 0.379 0.401
336 0.354 0.390 0.323 0.376 0.326 0.374 0.682 0.528 0.421 0.425
720 0.433 0.445 0.409 0.435 0.454 0.452 0.722 0.565 0.476 0.471
Avg. 0.345 0.381 0.322 0.371 0.329 0.373 0.607 0.506 0.403 0.417
ETTm2 96 0.172 0.265 0.169 0.259 0.183 0.273 0.223 0.295 0.208 0.283
192 0.228 0.306 0.223 0.295 0.223 0.301 0.281 0.329 0.344 0.372
336 0.281 0.345 0.293 0.341 0.278 0.339 0.364 0.373 0.354 0.383
720 0.403 0.424 0.451 0.433 0.425 0.424 0.475 0.435 0.460 0.455
Avg. 0.271 0.335 0.284 0.332 0.277 0.334 0.336 0.358 0.342 0.373
Weather 96 0.151 0.203 0.149 0.201 0.154 0.208 0.186 0.231 0.165 0.224
192 0.195 0.246 0.192 0.244 0.202 0.251 0.240 0.275 0.209 0.269
336 0.247 0.288 0.245 0.285 0.252 0.287 0.302 0.317 0.261 0.310
720 0.352 0.366 0.352 0.365 0.392 0.376 0.388 0.369 0.336 0.362
Avg. 0.236 0.275 0.234 0.273 0.250 0.280 0.279 0.298 0.243 0.291
Global Temp 96 0.192 0.328 0.192 0.329 0.189 0.322 0.260 0.390 0.210 0.344
192 0.238 0.375 0.236 0.375 0.234 0.376 0.301 0.423 0.253 0.385
336 0.259 0.397 0.256 0.397 0.253 0.399 0.359 0.464 0.282 0.411
720 0.345 0.465 0.322 0.451 0.292 0.426 0.371 0.477 0.342 0.457
Avg. 0.258 0.391 0.251 0.388 0.242 0.380 0.323 0.439 0.272 0.399
Average
0.306 0.362 0.304 0.359 0.301 0.358 0.416 0.422 0.364 0.400
@@ Appendix E Forecast Showcases
To visualize the performance differences among various time series foundation models, we present the forecasting results of our model, Time-MoE, in comparison to the ground truth across six real-world benchmarks. These benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Global Temp datasets. Alongside Time-MoE’s results, we also show the performance of other foundation models at different scales, providing a comprehensive view of their comparative capabilities (Figures 611). In all figures, the context length is set to 512, and the forecast horizon is 96. To enhance clarity and aesthetics, we display the full forecast output, complemented by a portion of the preceding historical input data, ensuring a more intuitive comparison.
@@
Figure 7: Zero-shot forecasting cases from ETTh2 by different models, with forecast horizon 96. Blue lines are the ground truths and read lines are the model predictions.
Figure 8: Zero-shot forecasting cases from ETTm1 by different models, with forecast horizon 96. Blue lines are the ground truths and read lines are the model predictions.
Figure 9: Zero-shot forecasting cases from ETTm2 by different models, with forecast horizon 96. Blue lines are the ground truths and read lines are the model predictions.
Figure 10: Zero-shot forecasting cases from Weather by different models, with forecast horizon 96. Blue lines are the ground truths and read lines are the model predictions.
Figure 11: Zero-shot forecasting cases from Global Temp by different models, with forecast horizon 96. Blue lines are the ground truths and read lines are the model predictions.